# A Critical Reanalysis of Supernova Type Ia Data

### v1

@@ 1 Introduction
Several attempts are underway to explain this tension. The explanations for the tension and possible resolution attempts include introduction of early dark energy models [33, 34] and also modified gravity theories [35, 36]. Recent findings by Riess et al. [37] show that combining JWST data yields estimates of , , and for Cepheid, JAGB, and TRGB methods, respectively. The combined result of aligns closely with HST estimates, suggesting that systematic biases in the distance ladder are unlikely to resolve the Hubble Tension. Whereas Friedman et al. [38] present new results from the Chicago Carnegie Hubble Program (CCHP) using JWST data, yielding values of for Tip of the Red-Giant Branch (TRGB) stars, for J-Branch Asymptotic Giant Branch (JAGB) stars, and for Cepheids. In [39], the authors propose an alternative approach to addressing the Hubble tension by focusing exclusively on "blue supernovae"—those with a colour parameter , which are less affected by dust extinction. By analysing only these supernovae in the Pantheon+ dataset, they derive a Hubble constant of , which is within of the Cosmic Microwave Background measurement. In contrast, using the full dataset yields a value of . This finding suggests that restricting the analysis to blue supernovae may help reconcile the Hubble tension.
@@ 4.1 Data Subsets
Employing the third binning scheme, we split the data in three (almost) sections with equal number of data points: , , and with redshift ranges: 0.0101 z 0.1769, 0.1771 z 0.3374 and 0.3375 z 2.26 respectively. The results of MCMC analysis for this are given in table 5). Comparing these with the full data set as we have done in the previous two binning schemes, we check the consistency of the full data set with individual bins (see fig 5) for both the likelihood functions. Here we can see that for logistic likelihood, the 2nd bin (pale green contour) raises the value of the Hubble constant for both the models. The higher redshift data, the 3rd bin (grey contour) is more consistent compared to other two bins for all the parameters.
@@ 6 Summary & Conclusion
In this work, we presented a detailed analysis of Pantheon compilation of supernova type Ia dataset, considering both the full dataset, binned data on the basis of redshift and the constituent surveys. We explored the deviation from Gaussian distribution of residuals in the distance modulus. Different parameterisations for equation of state parameter of dark energy were considered and various statistical goodness of fit tests verified that Logistic likelihood function fits better to the Pantheon data compared to the Gaussian likelihood. This is true for the full dataset and lower redshift bins whereas for the higher redshift bins, the comparison is inconclusive. The Markov Chain Monte Carlo method for parameter estimation using the Logistic likelihood for the full dataset suggests a value of the Hubble constant closer to that obtained from late-time measurements with larger error bars. Therefore choice of statistics can provide a way in reducing Hubble tension, in this case at the cost of increased uncertainties. This holds true irrespective of the dark energy parametrisations chosen for the analysis. In the data subsets analysis, it has been noted that higher redshift bins yield lower values of the Hubble constant when compared to lower redshift bins. This observation is consistent across the cosmological constant model. Our results for the cosmological constant are consistent with the ones obtained in [40] and we have extended the analysis to dynamical dark energy models and different subsets of the data.

### v2

@@ 1 Introduction
Several attempts are underway to explain this tension. The explanations for the tension and possible resolution attempts include introduction of early dark energy models [33, 34], modified gravity theories [35, 36, 37], Interaction between Dark Matter and Dark Energy [38], etc. Recent findings by Riess et al. [39] show that combining JWST data yields estimates of , , and for Cepheid, JAGB, and TRGB methods, respectively. The combined result of aligns closely with HST estimates, suggesting that systematic biases in the distance ladder are unlikely to resolve the Hubble Tension. Whereas Friedman et al. [40] present new results from the Chicago Carnegie Hubble Program (CCHP) using JWST data, yielding values of for Tip of the Red-Giant Branch (TRGB) stars, for J-Branch Asymptotic Giant Branch (JAGB) stars, and for Cepheids. In [41], the authors propose an alternative approach to addressing the Hubble tension by focusing exclusively on "blue supernovae"—those with a colour parameter , which are less affected by dust extinction. By analysing only these supernovae in the Pantheon+ dataset, they derive a Hubble constant of , which is within of the Cosmic Microwave Background measurement. In contrast, using the full dataset yields a value of . This finding suggests that restricting the analysis to blue supernovae may help reconcile the Hubble tension.
@@ 4.1 Data Subsets
Employing the third binning scheme, we split the data in three (almost) sections with equal numbers of data points: , , and with redshift ranges: 0.0101 z 0.1769, 0.1771 z 0.3374 and 0.3375 z 2.26 respectively. The results of MCMC analysis for this are given in table 5). It has been suggested that a binned analysis of Pantheon data shows that Hubble constant scales with redshift in the form of power-law as depicted in [78, 79] Comparing these with the full data set as we have done in the previous two binning schemes, we check the consistency of the full data set with individual bins (see fig 5) for both the likelihood functions. Here we can see that for logistic likelihood, the 2nd bin (pale green contour) raises the value of the Hubble constant for both the models. The higher redshift data, the 3rd bin (grey contour) is more consistent compared to other two bins for all the parameters.
@@ 6 Summary & Conclusion
In this work, we presented a detailed analysis of Pantheon compilation of supernova type Ia dataset, considering both the full dataset, binned data on the basis of redshift and the constituent surveys. We explored the deviation from Gaussian distribution of residuals in the distance modulus. Different parameterisations for equation of state parameter of dark energy were considered and various statistical goodness of fit tests verified that Logistic likelihood function fits better to the Pantheon data compared to the Gaussian likelihood. This is true for the full dataset and lower redshift bins whereas for the higher redshift bins, the comparison is inconclusive. The Markov Chain Monte Carlo method for parameter estimation using the Logistic likelihood for the full dataset suggests a value of the Hubble constant closer to that obtained from late-time measurements with larger error bars. Therefore choice of statistics can provide a way in reducing Hubble tension, in this case at the cost of increased uncertainties. This holds true irrespective of the dark energy parametrisations chosen for the analysis. In the data subsets analysis, it has been noted that higher redshift bins yield lower values of the Hubble constant when compared to lower redshift bins. This result is consistent with the results from [84, 85, 86]. This observation is consistent across the cosmological constant model. Our results for the cosmological constant are consistent with the ones obtained in [42] and we have extended the analysis to dynamical dark energy models and different subsets of the data.